OpenStudy (anonymous):

Hey guys, I'm trying to figure out how to "algebraically" prove there are no x-intercepts for graph of y = 1+sin 2x?... help please

OpenStudy (anonymous):

it has x-intercepts

OpenStudy (anonymous):

what ? how?... isn't it moved up 1 on the y axis with an amplitude of 1?

zepdrix (zepdrix):

So instead of going up to 1 and down to -1, it will go up to 2 and down to 0 (intercepting/touching the axis) yes?

OpenStudy (anonymous):

You lose.

OpenStudy (anonymous):

correct. but it's not necessarily intersecting right?

zepdrix (zepdrix):

It `intercepts` the x-axis. But yes you are correct it doesn't necessarily pass through which I assume is what you mean by `intersect`.

OpenStudy (anonymous):

so if I set y= o and then solve for x i would get 0= 1+sin2x -1 = sin 2x sin = -1 @ pi/2 & 3pi/2 , so, pi/2 = 2x pi/4 =x

zepdrix (zepdrix):

Only -1 at 3pi/2. 3pi/2 = 2x

OpenStudy (anonymous):

oh,yea duh. so the only "intercept" is 3pi/4. correct?

zepdrix (zepdrix):

No there is an infinite number of intercepts since sin2x is periodic. I guess a better way to word it would be, sin2x = -1 implies,\[\Large\bf\sf 2x\quad=\quad \frac{3\pi}{2}+2k \pi\]

OpenStudy (anonymous):

no, there's actually infinitely many x-intercept

zepdrix (zepdrix):

Where k is an integer, adding or subtracting full rotations gives us another intercept.

zepdrix (zepdrix):

It's important to write the 2kpi BEFORE dividing by 2, because it affects the other solutions.

zepdrix (zepdrix):

\[\Large\bf\sf x\quad=\quad \frac{3\pi}{4}+k \pi\]

zepdrix (zepdrix):

When k=0, we get the expected solution 3pi/4. When k=1, we get 7pi/4. When k=2, we get 11pi/4. And so on...

zepdrix (zepdrix):

Confused by any of that? :o

OpenStudy (anonymous):

I'm kind of getting confused, so I have to solve setting the 2x equal to the 3pi first? or thats only for the finding all other infinite solutions?

zepdrix (zepdrix):

Setting 2x equal to 3pi/2 only finds the first solution. Setting 2x equal to 3pi/2 + 2pi is the second solution. Understand how those two solutions would fall on the same spot on the unit circle but the second one has rotated around a full time?

zepdrix (zepdrix):